The sum of the ages of Jones' four children is 25. Maria's oldest brother is six times as old as she is. Her next-to-oldest brother is 5 years older than she is, and the youngest brother is twice Maria's age. What are the ages of the Jones' four children?
Answer by ankor@dixie-net.com(22740)

The sum of the ages of Jones' four children is 25. Maria's oldest brother is six times as old as she is.
Her next-to-oldest brother is 5 years older than she is, and the youngest brother is twice Maria's age.
What are the ages of the Jones' four children?
:
Let m = M's age
:
Write an equation for each statement
:
"Maria's oldest brother is six times as old as she is.
6m = oldest brother
:
"Her next-to-oldest brother is 5 years older than she is,"
m+5 = next brother
;
" the youngest brother is twice Maria's age."
2m = youngest brother
:
"The sum of the ages of Jones' four children is 25.
6m + (m+5) + 2m + m = 25
10m = 25 - 5
10m = 20
m = 2 yrs is M's age
:
You can find the ages of the brothers from this